Write an equation in standard form of the line that passes through (2,-3) and (-3,7).
Answer by rfer(16322)

m=(7--3)/(-3-2)
m=10/-5
m=-2
------------------
y=mx+b
7=-2*-3+b
7=6+b
b=1
=============
-2x+y=1
2x-y=-1